Elemental analysis of 5.00 g compound shows that it contained 1.278 g of carbon, 0.318 g of hydrogen and 3.404 g of sulfur. The molecular mass of the compound was determined in a experiment of this compound.
Answer #1

One way of doing this problem is to use percentages of the elements in the compound and compare them to the elements in a mole of the compound

total mass of elements in the sample = 1.278+.318+3.40 = 5.0g

% C= 1.278 / 5 = 25.56%
%H = .318 / 5.0 = 6.36%
%S = 3.40 / 5.0 = 68.0 %

If we now use these percentages to work out the mass of each element in a mole we get

mass of C = 25.56 % of 94.19g = 24 g or 24 / 2 moles of C = 2
mass of H = 6.36 % of 94.19g = 6 g of H or 6 moles of H
mass of S = 68 % of 94.19g = 64g or 2 moles of S

molecular formula is C2H6S2 ...